# IIT Gandhinagar Gujarat, India Arxiv

### v3

@@ Abstract
Neural radiance fields (NeRF) have revolutionized photorealistic rendering of novel views for 3D scenes. Despite their growing popularity and efficiency as 3D resources, NeRFs face scalability challenges due to the need for separate models per scene and the cumulative increase in training time for multiple scenes. The potential for incrementally encoding multiple 3D scenes into a single NeRF model remains largely unexplored. To address this, we introduce Continual-Neural Graphics Primitives (C-NGP), a novel continual learning framework that integrates multiple scenes incrementally into a single neural radiance field. Using a generative replay approach, C-NGP adapts to new scenes without requiring access to old data. We demonstrate that C-NGP can accommodate multiple scenes without increasing the parameter count, producing high-quality novel-view renderings on synthetic and real datasets. Notably, C-NGP models all $8$ scenes from the Real-LLFF dataset together, with only a $2.2\%$ drop in PSNR compared to vanilla NeRF, which models each scene independently. Further, C-NGP allows multiple style edits in the same network.
@@ 3.1 Background
Neural Radiance Fields. Mildenhall et al. [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] introduced NeRF, a method for synthesizing novel views from a set of input images using a multi-layer perceptron (MLP) that maps a 5D input—3D spatial coordinates $(x,y,z)$ and viewing direction $(\theta,\phi)$—to the density $\sigma$ and color $c$ of a point. Points are sampled along camera rays $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ is the camera origin, $\mathbf{d}$ is the ray direction, and $t$ is the sampled depth. Rendering is performed via volumetric integration [Max(1995)], computing the color of a ray as $\hat{C}(r)=\sum_{i=1}^{N}w_{i}c_{i}$, where the weights $w_{i}=T_{i}\alpha_{i}$ depend on transmittance $T_{i}$ and opacity $\alpha_{i}=(1-e^{-\sigma_{i}\delta_{i}})$, with $\delta_{i}$ being the distance between adjacent samples. The model is trained by minimizing the squared error between predicted and ground-truth pixel colors: $\mathcal{L}=\sum_{r\in\mathcal{R}}||\hat{C}(r)-C(r)||_{2}^{2}$, where $\mathcal{R}$ denotes the set of rays. Inputs to the MLP are encoded using sinusoidal positional encodings to better capture high-frequency scene details.
@@ 4.2 Training Details
For training C-NGP, we used a batch size of $10,000$ rays and an initial learning rate of $2\times 10^{3}$. The model was trained for $30$ epochs across synthetic and real datasets. C-NGP was conditioned on pseudo labels assigned to each scene to ensure scene-specific representations. The hash table size for Instant-NGP was set to $T=2^{19}$, with a feature size of $F=4$. For the NeRF Synthetic $360^{\circ}$ dataset, we applied the generative replay method [Shin et al.(2017)Shin, Lee, Kim, and Kim], where previously observed scenes were re-rendered using the camera parameters of the new scene. For real-world datasets like LLFF [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] and Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], which lack full $360^{\circ}$ coverage, we only store the camera parameters of each scene. This allowed us to generate novel views while avoiding artifacts such as floaters or ghosting, less problematic issues in synthetic datasets with complete $360^{\circ}$ coverage.
@@ 4.3 Experimental Evaluation
Method Current Scene [L] Current Scene [L $\xrightarrow{}$ C] Current Scene [L $\xrightarrow{}$ C $\xrightarrow{}$ S] Current Scene [L $\xrightarrow{}$ C $\xrightarrow{}$ S $\xrightarrow{}$ D]
@@ 4.4 Additional Results and Analysis
Table 3: (Left) Analysing upper bound on the number of scenes C-NGP can accommodate with reasonable loss in rendering quality across single and mixed datasets. Single: $22$ scenes from BlenderNeRF dataset. Mixed: 14 scenes from BlenderNeRF + $8$ scenes from NeRF Synthetic $360^{\circ}$ dataset. (Right) Ablation on the configuration choice for conditioning Instant-NGP network [Müller et al.(2022)Müller, Evans, Schied, and Keller] - the backbone of C-NGP.
@@
Supplementary. Due to space constraints, we elaborate: (a) quantitative comparison with other methods, (b) time analysis, (c) shared feature space visualization, and (d) additional discussion in the supplementary material. We urge the readers to kindly refer to the supplementary material as well for a holistic understanding of this work.
@@ 5 Conclusion
We introduced C-NGP, a continual learning framework for incremental multi-scene modeling within a single neural radiance field. C-NGP accommodates new scenes while reasonably preserving the previously learned information via scene conditioning and generative replay. Extensive empirical studies on synthetic and real datasets demonstrate that C-NGP achieves strong performance in multi-scene representation learning and outperforms existing methods over training and inference speed, offering real-time rendering capabilities. While at this stage, C-NGP incurs an acceptable loss in rendering quality while modeling multiple scenes, its performance in highly diverse scene settings and attempt towards zero forgetting warrants further exploration.
@@ Appendix A Quantitative Comparison with Other Methods
Method Per Scene [S, F, H, L] Fine-tune [S $\xrightarrow{}$ F] Fine-tune [S $\xrightarrow{}$ F $\xrightarrow{}$ H] Fine-tune [S $\xrightarrow{}$ F $\xrightarrow{}$ H $\xrightarrow{}$ L]
@@ Appendix C Shared Representation Space
/* Render previous scenes and store in $I_{prev}$/
@@
       $I_{prev}$.append($render\_image$)

### v4

# IIT Gandhinagar Gujarat, India Incremental Multi-Scene Modeling via C-NGP

@@ Abstract
Neural radiance fields (NeRF) have revolutionized photorealistic rendering of novel views for 3D scenes. Despite their growing popularity and efficiency as 3D resources, NeRFs face scalability challenges due to the need for separate models per scene and the cumulative increase in training time for multiple scenes. The potential for incrementally encoding multiple 3D scenes into a single NeRF model remains largely unexplored. To address this, we introduce Continual-Neural Graphics Primitives (C-NGP), a novel continual learning framework that integrates multiple scenes incrementally into a single neural radiance field. Using a generative replay approach, C-NGP adapts to new scenes without requiring access to old data. We demonstrate that C-NGP can accommodate multiple scenes without increasing the parameter count, producing high-quality novel-view renderings on synthetic and real datasets. Notably, C-NGP models all $8$ scenes from the Real-LLFF dataset together, with only a $2.2\%$ drop in PSNR compared to vanilla NeRF, which models each scene independently. Further, C-NGP allows multiple style edits in the same network. The code implementation and dynamic visualizations can be accessed from here https://prajwalsingh.github.io/C-NGP/.
@@ 3.1 Background
[…]
@@ 4.2 Training Details
For training C-NGP, we used a batch size of $10,000$ rays and an initial learning rate of $2\times 10^{-3}$. The model was trained for $30$ epochs across synthetic and real datasets. C-NGP was conditioned on pseudo labels assigned to each scene to ensure scene-specific representations. The hash table size for Instant-NGP was set to $T=2^{19}$, with a feature size of $F=4$. For the NeRF Synthetic $360^{\circ}$ dataset, we applied the generative replay method [Shin et al.(2017)Shin, Lee, Kim, and Kim], where previously observed scenes were re-rendered using the camera parameters of the new scene. For real-world datasets like LLFF [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] and Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], which lack full $360^{\circ}$ coverage, we only store the camera parameters of each scene. This allowed us to generate novel views while avoiding artifacts such as floaters or ghosting, less problematic issues in synthetic datasets with complete $360^{\circ}$ coverage.
@@ 4.3 Experimental Evaluation
While C-NGP shows strong performance on synthetic datasets and Tanks & Temples, its performance on unbounded LLFF scenes is relatively lower. We attribute this partly to the Instant-NGP backbone, which is known to be less effective in such settings [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman, Li et al.(2023)Li, Gao, Tancik, and Kanazawa]. Nonetheless, C-NGP shows better performance than Instant-NGP after fine-tuning while retaining efficiency and scalability.
@@
[…]
@@ 4.4 Additional Results and Analysis
Table 3: (Left) Analysing upper bound on the number of scenes C-NGP can accommodate with reasonable loss in rendering quality across single and mixed datasets. Single: $22$ scenes from BlenderNeRF dataset. Mixed: 14 scenes from BlenderNeRF + 88 scenes from NeRF Synthetic $360^{\circ}$ dataset. (Right) Ablation on the configuration choice for conditioning Instant-NGP network [Müller et al.(2022)Müller, Evans, Schied, and Keller] - the backbone of C-NGP.
@@
Supplementary. Due to space constraints, we elaborate: (a) quantitative comparison with other methods, (b) time analysis, (c) shared feature space visualization, and (d) additional discussion in the supplementary material. We urge the readers to kindly refer to the supplementary material as well for a holistic understanding of this work. These additional studies complement the main results and ensure that our evaluation is comprehensive, addressing performance, efficiency, and representation quality.
@@ 5 Conclusion
We introduced C-NGP, a continual learning framework for incremental multi-scene modeling within a single neural radiance field. The proposed method accommodates new scenes while reasonably preserving the previously learned information via scene conditioning and generative replay. Extensive empirical studies on synthetic and real datasets demonstrate that our method achieves strong performance in multi-scene representation learning and outperforms existing methods over training and inference speed, offering real-time rendering capabilities. C-NGP offers practical value for scalable 3D repositories, enabling multiple scenes and even style edits to be stored and shared in a single lightweight checkpoint, avoiding the overhead of per-scene models. While at this stage, the method incurs an acceptable loss in rendering quality while modeling multiple scenes, its performance in highly diverse scene settings and attempt towards zero forgetting warrants further exploration.
6 Acknowledgment
This work was supported by the Prime Minister Research Fellowship awarded to Prajwal Singh and Ashish Tiwari, and by the Jibaben Patel Chair in Artificial Intelligence held by Shanmuganathan Raman.
@@ Appendix A Quantitative Comparison with Other Methods
[…]
@@ Appendix C Shared Representation Space
/* Render previous scenes and store in $I_{prev}$ */
@@
 $I_{prev}$.append(render_imagerender\_image)